OpenStudy (anonymous):

FAN AND MEDAL!! Heinz has a list of possible functions. Pick one of the g(x) functions below, show how to find the zeros, and then describe to Heinz the other key features of g(x). g(x) = x^3 – x^2 – 4x + 4 g(x) = x^3 + 2x^2 – 9x – 18 g(x) = x^3 – 3x^2 – 4x + 12 g(x) = x^3 + 2x^2 – 25x – 50 g(x) = 2x^3 + 14x^2 – 2x – 14

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@Cosmichaotic

OpenStudy (anonymous):

Which polynomial did you do? I had the option to complete one of the five I mentioned.

OpenStudy (anonymous):

We find the zeroesby setting y = 0 or g(x) = 0, since g(x) = y and then seeing what would cause it to be 0. The solutions (We also call them roots or zeroes), are 4, -1, and 0 Because if we were to substitute any of these values in for x, we would get 0.

OpenStudy (anonymous):

I just did the first one.

OpenStudy (anonymous):

g(x) = x^3 – x^2 – 4x + 4

OpenStudy (anonymous):

Wait a sec, I factored that crap wrong heh

OpenStudy (anonymous):

(x+2)(x-2)(x+1)

OpenStudy (anonymous):

So the zeroes are x = -2, 2, and -1 That is where the graph crosses the x-axis in the graph.

OpenStudy (anonymous):

So the problem asks me to show how to find the zeros. Even if its not completely accurate, can you provide me with a basic rundown of what I can show to get me by? I have to complete 12 assignments in a row by tomorrow morning at 8. Haha

OpenStudy (anonymous):

Yea. Um, we find the zeroes by setting the polynomial equal to 0 and factoring it. Then we see what values of x will make the polynomial 0, and those are our zeroes.

OpenStudy (anonymous):

Like if our polynomial is x^2+2x+1, then if we set it to 0 we get x^2+2x+1=0 Then if we factor we get (x+1)(x+1)=0. And if we set x = -1 then this will come out to be 0 and will work. So x = -1 is a zero of the polynomial x^2+2x+1... that's just an example hehe

OpenStudy (anonymous):

I can explain better. I will tomorrow after I do a little more studying myself. I need to figure out how to factor higher degree polynomials and finding their 0's tonight.

OpenStudy (anonymous):

Okay thank you!